# Syntax-Guided Procedural Synthesis of Molecules

### v1

@@ Appendix H Docking Case Study with AutoDock Vina
(b) Top sample binders against MPro from literature, based on consensus docking scores [23]
In this section, we structurally analyze the top molecules discovered by our method, visualized in 15(a).
For our optimized binders against DRD3, the chlorine substituent and polycyclic aromatic structure suggest good potential for binding through $\pi-\pi$ interactions and halogen bonding. The bromine and carboxyl groups can enhance binding affinity through halogen bonding and hydrogen bonding, respectively. The polycyclic structure further supports $\pi-\pi$ stacking interactions. In general, they have a comparable binding capability than the baseline molecules, but with simpler structures, the ease of synthesis for the predicted molecules are higher than the baseline molecules.
For our optimized binders against Mpro, the three predicted molecules contain multiple aromatic rings in conjugation with halide groups. The conformation structures of the multiple aligned aromatic rings play a significant role in docking and achieve ideal molecular pose and binding affinity to Mpro, compared with the baseline molecules shown in 15(b). The predicted structures also indicate stronger $\pi-\pi$ interaction and halogen bonding compared with the baselines. In terms of ease of synthesis, Bromination reactions are typically straightforward, but multiple fused aromatic rings can take several steps to achieve. In general, the second and third can be easier to synthesize than Brc1cc(-c2cc(-c3cccc4ccccc34)nc3ccccc23)c2ccccc2n1 due to less aromatic rings performed. However, the literature molecules appeared to be even harder to synthesize due to their high complexicity structures. So the predicted molecules obtained a general higher ease of synthesis than the baseline molecules. Compared with the other baseline molecules, e.g. Manidipine, Lercanidipine, Efonidipine (Dihydropyridines), known for their calcium channel blocking activity, but not specifically protease inhibitors, Azelastine, Cinnoxicam, Idarubicin vary widely in their primary activities, not specifically designed for protease inhibition. Talampicillin and Lapatinib are also primarily designed for other mechanisms of action. Boceprevir, Nelfinavir, Indinavir, on the other hand, are known protease inhibitors with structures optimized for binding to protease active sites, so can serve as strong benchmarks. Overall, the binding effectiveness of the predicted molecules are quite comparable to the baseline molecules.

### v2

@@ Appendix I Docking Case Study with AutoDock Vina
(b) Top binders against Mpro from literature, based on consensus docking scores (Ghahremanpour et al., 2020)
We structurally analyze the top molecules discovered by our method, visualized in Figure 18(a).
For our optimized binders against DRD3, the chlorine substituent and polycyclic aromatic structure suggest good potential for binding through $\pi-\pi$ interactions and halogen bonding. The bromine and carboxyl groups can enhance binding affinity through halogen bonding and hydrogen bonding, respectively. The polycyclic structure further supports $\pi-\pi$ stacking interactions. In general, they have a comparable binding capability to the baseline molecules, but with simpler structures, so the ease of synthesis for the predicted molecules are higher than the baseline molecules.
For our optimized binders against Mpro, the three predicted molecules contain multiple aromatic rings in conjugation with halide groups. The conformation structures of the multiple aligned aromatic rings play a significant role in docking and achieve ideal molecular pose and binding affinity to Mpro, compared to the baseline molecules shown in Figure 18(b). The predicted structures indicate stronger $\pi-\pi$ interaction and halogen bonding compared with the baselines. In terms of ease of synthesis, Bromination reactions are typically straightforward, but multiple fused aromatic rings can take several steps to achieve. In general, the second and third can be easier to synthesize than the top binder due to less aromatic rings performed. However, the literature molecules appeared to be even harder to synthesize due to their high complexity structures. So the predicted molecules obtained a general higher ease of synthesis than the baseline molecules. Compared with the other baseline molecules, e.g. Manidipine, Lercanidipine, Efonidipine (Dihydropyridines), known for their calcium channel blocking activity, but not specifically protease inhibitors, Azelastine, Cinnoxicam, Idarubicin vary widely in their primary activities, not specifically designed for protease inhibition. Talampicillin and Lapatinib are also primarily designed for other mechanisms of action. Boceprevir, Nelfinavir, Indinavir, on the other hand, are known protease inhibitors with structures optimized for binding to protease active sites, so can serve as strong benchmarks. Overall, the binding effectiveness of the predicted molecules are quite comparable to the baseline molecules.